### v1

@@ 4 Numerical Experiments
In this section, we discuss the numerical performance of our proposed method (MDDLSCG), comparing it with Algorithm 3.1 in [26] (denoted by MSCG) and Scaled conjugate gradient method (denoted by ScCGM) in [29]. All numerical experiments were conducted in MATLAB R2024b on an HP workstation with Intel Core i7-13700 processor, 32 GB RAM and Window 11.
During the numerical experiments, we set the parameters $\delta=0.01$, $\sigma=0.1$ for all the algorithms, $\eta=0.001$, $\tau=10$, $r=1$ and $\nu=0.001$ for algorithms MSCG and MDDLSCG and $p=0.4$ and $q=0.2$ for the algorithm MDDLSCG. During the implementation, the iterations are terminated, when it meets the stopping criterion $E_{n}=\|\nabla f(x_{n})\|_{\infty}<10^{-15}$.
For the numerical experiment, we consider the test problem of finding the minima of Beale function. The Beale function is multimodal with sharp peaks at the corners of the input domain, which is defined from the general Beale function as follows:
@@
The numerical results of all the algorithms (MDDLSCG, MSCG and ScCG) are summarized in Table 1, while their performance is illustrated in Figure LABEL:bealee. In Figure LABEL:bealee (a), the iteration paths are visualized on the contour plot of the Beale function, showing how each method converges to the minimizer. Figure LABEL:bealee (b) depicts the convergence rate by plotting the norm of the gradient, $\|\nabla f(x_{n})\|_{\infty}$ versus the number of iterations on a logarithmic scale.
From the data presented, we can clearly observe that our method, MDDLSCG, outperforms the MSCG algorithm and ScCG in all tested scenarios. This superiority is evidenced by the metrics shown in the table and figures, which highlight the efficiency and effectiveness of MDDLSCG compared to MSCG and ScCG.
Method itr Tcpu $E_{n}$
MDDLSCG 21 5.957900e-03 3.580469e-15
MSCG 30 1.783450e-02 4.662937e-15
ScCG 139 9.749100e-03 8.840151e-15
Table 1: Comparison table for MDDLSCG, MSCG and ScCG for Beale problems
@@ 5 Application in compressed sensing
where $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is a penalty function, and $\mu>0$ is a parameter that controls the trade-off between sparsity and reconstruction fidelity. A popular choice for $\phi$ is $\phi(x)=\|x\|_{1}$, leading to the well-known Basis Pursuit Denoising (BPD) problem, which has been extensively studied in the literature; see [32] and the references therein. However, solving BPD is challenging due to the non-smoothness of the $\ell_{1}$-penalty term. To address this, Zhu et al. [33] proposed a relaxation of BPD using Nesterov’s smoothing technique [34], resulting in the following formulation:
@@
$\psi_{\lambda}(|x_{i}|)=\max_{\gamma\in[0,1]}\left\{\gamma|x_{i}|-\frac{1}{2}\lambda\gamma^{2}\right\}$
@@
For the numerical experiments, we aim to reconstruct a sparse signal of length $n$ from $m$-length observations ($m\ll n$). The parameters for the MDDLSCG, MSCG and ScCG algorithms are set as follows: $\mu=\max\{2^{-7},0.001\|A^{T}b\|_{\infty}\}$, $\lambda=\min\{0.001,0.048\|A^{T}b\|_{\infty}\}$, $\sigma=0.1$, $\delta=0.01$, $\eta=0.001$, $\tau=10$, $r=1$ and $\nu=0.001$.
@@
To evaluate the restoration performance qualitatively, we report the relative error (RelErr) of the recovered signal, defined as:
@@
The detailed numerical results are presented in Table 2, where we report the average number of iterations (Itr), average CPU time (Tcpu), mean squared error (MSE), and relative error (RelErr) for each setting. The graph depicting the relationship between relative error and the number of iterations is shown in Figures LABEL:stpp512(a) and LABEL:stopp(a), while the graph illustrating MSE versus the number of iterations is presented in Figures LABEL:stpp512(b) and LABEL:stopp(b). Additionally, the recovery of the signal is visualized in Figures 4 and 6.
@@
Dimensions $(m,n,k)$ Method Itr Tcpu MSE RelErr
(128,512,16) MDDLSCG 272 1.186382e-01 2.033486e-06 7.487872e-03
MSCG 493 1.606987e-01 2.033495e-06 7.488087e-03
ScCG 3653 1.135658e+00 2.033588e-06 7.712389e-03
(256,1024,32) MDDLSCG 291 2.163503e-01 2.128598e-06 9.721181e-03
MSCG 1405 8.339790e-01 2.128516e-06 9.720993e-03
ScCG 3002 1.746555e+00 2.229608e-04 9.949161e-02
@@ 6 Conclusions
We presented a modified descent Dai-Liao spectral conjugate gradient (MDDLSCG) method that incorporates a new secant condition and quasi-Newton direction. The updated method introduces spectral parameters ensuring a sufficient descent property independent of line search techniques. Theoretical analysis establishes the global convergence of the method for general nonlinear functions under standard assumptions. Numerical experiments and an application in signal processing demonstrate the improved performance and practical effectiveness of the proposed algorithm, highlighting its potential in optimization and related applications.

### v2

@@ 4 Numerical Experiments
In this section, we discuss the numerical performance of our proposed method (MDDLSCG), comparing it with Algorithm 3.1 in [26] (denoted by MSCG) and Scaled conjugate gradient method (denoted by ScCG) in [28]. All numerical experiments were conducted in MATLAB R2024b on an HP workstation with Intel Core i7-13700 processor, 32 GB RAM and Windows 11.
For the numerical experiment, we consider two test problems: (i) finding the minima of the Beale function, and (ii) finding the minima of the Almost Perturbed Quadratic (APQ) function:
Beale Test Function: The Beale function is multimodal with sharp peaks at the corners of the input domain, which is defined from the general Beale function as follows:
@@
During the numerical experiments, we set the parameters $\delta=0.01$, $\sigma\in\{0.1,0.2,0.4,0.6,0.8,0.9\}$ for all the algorithms, $\eta=0.001$, $\tau=10$, $r=1$ and $\nu=0.001$ for algorithms MSCG and MDDLSCG and $p=0.4$ and $q=0.2$ for the algorithm MDDLSCG. We take the initial point $x_{0}=(1,0.8)$. During the implementation, the iterations are terminated, when it meets the stopping criterion $E_{n}=\|\nabla f(x_{n})\|_{\infty}<10^{-15}$.
@@
The numerical results of all the algorithms MDDLSCG, MSCG and ScCG for the Beale function are summarized in Table 1, while their performance is illustrated in Figure LABEL:bealefig1 and LABEL:bealefig2. In Figures LABEL:bealefig1 (a), (c), (e), the iteration paths are visualized on the contour plot of the Beale function, showing how each method converges to the minimizer for $\sigma=0.1,0.2,0.6$, respectively. Figures LABEL:bealefig1 (b), (d), (f) depict the convergence rate by plotting the norm of the gradient, $\|\nabla f(x_{n})\|_{\infty}$ versus the number of iterations on a logarithmic scale for $\sigma=0.1,0.2,0.6$, respectively. The graph between different values of sigma and number of iterations is given in Figure LABEL:bealefig2 (a) and the graph between different values of sigma and execution time is in Figure LABEL:bealefig2 (b).
From the data presented, it is evident that our proposed method, MDDLSCG, outperforms both MSCG and ScCG in all tested scenarios. As shown in Table 1, Figure LABEL:bealefig1 and LABEL:bealefig2 (b), MDDLSCG reaches the stopping criteria in fewer iterations and requires less execution time. Additionally, From Figure LABEL:bealefig2 (a), we observe that these methods are significantly sensitive to variations in the parameter $\sigma$, which affects their performance.
$\sigma$ MDDLSCG MSCG ScCG itr Tcpu $E_{n}$ itr Tcpu $E_{n}$ itr Tcpu $E_{n}$ 0.1 37 1.86e-02 9.49e-15 63 3.14e-02 2.63e-16 200 5.40e-02 5.27e-16 0.2 36 1.75e-02 3.16e-15 63 2.76e-02 9.07e-15 136 3.99e-02 4.51e-15 0.4 53 2.44e-02 9.40e-15 73 2.65e-02 8.75e-15 126 4.73e-02 2.87e-15 0.6 40 2.36e-02 7.91e-16 72 2.78e-02 5.74e-15 95 4.94e-02 9.32e-15 0.8 67 3.24e-02 5.32e-15 111 2.86e-02 4.48e-15 201 4.92e-02 6.91e-15 0.9 64 3.28e-02 8.16e-15 101 3.48e-02 8.74e-15 184 4.37e-02 1.08e-15
Table 1: Comparison table for MDDLSCG, MSCG and ScCG for Beale function
Almost Perturbed Quadratic function: Consider the Almost Perturbed Quadratic (APQ) function $f:\mathbb{R}^{m}\to\mathbb{R}$ defined by
$f(\mathbf{x})=\frac{1}{2}\mathbf{x}^{T}A\mathbf{x}+\mathbf{b}^{T}\mathbf{x}$
where $A\in\mathbb{R}^{m\times m}$ is a symmetric positive definite matrix ans $b\in\mathbb{R}^{m}$ is a vector of perturbation. This function has a global minimum at $x=-A^{-1}b$. Note that the gradient of $f$ is given by
$\nabla f(x)=Ax+b\text{ for all }x\in\mathbb{R}^{m}.$
For the numerical experiments, we have considered three cases, each characterized by dimensions of1000100010001000, $100$ and $25$. To the cases with dimensions $1000$ and $100$, the matrix $A$ and vector $b$ are randomly generated using MATLAB’s built-in function rand and the initial point is randomly generated using MATLAB’s rand function.
For the case with dimension $25$, we take the initial point
$x_{0}$ $=(x_{0,1},\ldots,x_{0,i},\ldots,x_{0,25}),\text{ where }x_{0,i}=1\text{ for all }i=1,2,\ldots,25\text{ and}$
$a$ $=(2,4,3,6,4,12,8,10,5,8,6,11,4,8,3,6,2,12,6,8,4,10,5,2,6),$
$b$ =(4,52,18,60,24,72,72,50,55,80,78,22,12=(-4,-52,-18,-60,-24,-72,-72,-50,-55,-80,-78,-22,-12,= ( - 4 , - 52 , - 18 , - 60 , - 24 , - 72 , - 72 , - 50 , - 55 , - 80 , - 78 , - 22 , - 12 ,
80,27,42,24,120,30,120,24,100,60,20,66)$\quad$-80,-27,-42,-24,-120,-30,-120,-24,-100,-60,-20,-66),- 80 , - 27 , - 42 , - 24 , - 120 , - 30 , - 120 , - 24 , - 100 , - 60 , - 20 , - 66 ) ,
$A$ $=\text{diag}(a).$
For all the cases, the parameters are set as follows: $\delta=0.01$ and $\sigma=0.1$ for all algorithms, while for MSCG and MDDLSCG, we use $\eta=0.001$, $\tau=10$, $r=1$, and $\nu=0.001$. Additionally, for the MDDLSCG algorithm, we set $p=0.4$ and $q=0.2$.
During the implementation, the iterations are terminated when the stopping criterion $E_{n}=\|\nabla f(x_{n})\|_{\infty}<10^{-6}$ is satisfied, where $x_{n}=(x_{n,1},\ldots,x_{n,25})$.
The performance of the algorithms MDDLSCG, MSCG, and ScCG for the APQ function is summarized in Table 2 for all three cases. To provide a clearer visualization of the experiment, we have plotted the graph of the number of iterations versus $E_{n}$ in Figure LABEL:APQgraph. From Table 2 and Figure LABEL:APQgraph, it is evident that our proposed method, MDDLSCG, requires fewer iterations and less computational time to satisfy the desired stopping criterion across all cases.
Figure LABEL:APQtrack illustrates how the algorithms MDDLSCG, MSCG, and ScCG approximate the desired solution $x^{*}=-A^{-1}b$ at the 5th and 10th iterations for the case where the problem dimension is $25$.
Furthermore, Figure LABEL:APQprogress depicts the convergence behavior of the algorithms as the number of iterations increases up to the 50th iteration for the same dimension. It is observed that MDDLSCG demonstrates superior convergence characteristics compared to MSCG and ScCG.
Dimension Method itr Tcpu $Er_{n}$
25 MDDLSCG 52 3.088055e-02 5.691882e-07
MSCG 59 3.947327e-02 6.931169e-07
ScCG 65 4.730704e-02 4.514777e-07
100 MDDLSCG 102 5.552326e-02 8.162825e-07
MSCG 121 6.512598e-02 9.411660e-07
ScCG 134 6.823947e-02 8.609522e-07
1000 MDDLSCG 132 3.481332e-02 9.291368e-07
MSCG 146 3.522395e-02 9.484382e-07
ScCG 171 4.053286e-02 8.747078e-07
Table 2: Comparison table of MDDLSCG, MSCG and ScCG for APQ function
@@ 5 Application in compressed sensing
where $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is a penalty function, and $\mu>0$ is a parameter that controls the trade-off between sparsity and reconstruction fidelity. A popular choice for $\phi$ is $\phi(x)=\|x\|_{1}$, leading to the well-known Basis Pursuit Denoising (BPD) problem, which has been extensively studied in the literature; see [32] and the references therein. However, solving BPD is challenging due to the non-smoothness of the $\ell_{1}$-penalty term. To address this issue, Zhu et al. [33] proposed a relaxation of BPD using Nesterov’s smoothing technique [34], resulting in the following formulation:
@@
$\psi_{\lambda}(|x_{i}|)=\max_{\gamma\in[0,1]}\left\{\gamma|x_{i}|-\frac{1}{2}\lambda\gamma^{2}\right\},$
@@
For the numerical experiments, we aim to reconstruct a sparse signal of length $n$ from $m$-length observations ($m\ll n$). The parameters for the MDDLSCG, MSCG and ScCG algorithms are set as follows:
$\mu=\max\{2^{-7},0.001\|A^{T}b\|_{\infty}\}$, $\lambda=\min\{0.001,0.048\|A^{T}b\|_{\infty}\}$, $\sigma=0.1$, $\delta=0.01$, $\eta=0.001$, $\tau=10$, $r=1$ and $\nu=0.001$.
@@
To evaluate the restoration performance qualitatively, we report the relative error (RelErr) of the recovered signal, which is defined as follows:
@@
The detailed numerical results are presented in Table 3, where we report the average number of iterations (Itr), average CPU time (Tcpu), mean squared error (MSE), and relative error (RelErr) for each setting. The graph depicting the relationship between relative error and the number of iterations is shown in Figures LABEL:stpp512(a) and LABEL:stopp(a), while the graph illustrating MSE versus the number of iterations is presented in Figures LABEL:stpp512(b) and LABEL:stopp(b). Additionally, the recovery of the signal is visualized in Figures 11 and 13.
@@
Dim$(m,n,k)$ Method Itr Tcpu MSE RelErr (128,512,16) MDDLSCG 272 1.186382e-01 2.033486e-06 7.487872e-03 MSCG 493 1.606987e-01 2.033495e-06 7.488087e-03 ScCG 3653 1.135658e+00 2.033588e-06 7.712389e-03 (256,1024,32) MDDLSCG 291 2.163503e-01 2.128598e-06 9.721181e-03 MSCG 1405 8.339790e-01 2.128516e-06 9.720993e-03 ScCG 3002 1.746555e+00 2.229608e-04 9.949161e-02
@@ 6 Conclusions
We investigated a modified descent Dai-Liao spectral conjugate gradient (MDDLSCG) method that incorporates a new secant condition and quasi-Newton direction. The updated method introduces spectral parameters ensuring a sufficient descent property independent of line search techniques. Theoretical analysis establishes the global convergence of the method for general nonlinear functions under standard assumptions. Numerical experiments and an application in signal processing demonstrate the improved performance and practical effectiveness of the proposed algorithm. Additionally, we present numerical results that validate its performance and show its superior convergence compared to the MSCG method [26] and the ScCG method [28], highlighting its potential in optimization and related applications.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
The authors are grateful to the editor and to the reviewers for their remarks and advice which improved our manuscript.
Funding
The authors declare that no funds, grants, or other support were received during the preparation of this manuscript.